OpenStudy (anonymous):

Solve the absolute value equation |2n - 4| = 8. One solution is n = 6. What is the other solution? I get no solution but what do u guys get?

OpenStudy (anonymous):

-8<=2n-4<=8 -4<=2n<=12 -2<=n<=6

OpenStudy (anonymous):

other solution -2

OpenStudy (hoblos):

|2n - 4| = 8 so 2n - 4= ±8 2n-4=8 or 2n-4=-8 n=6 n=-2